OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 36x2 − 18x − 9 4a2 − 28a − 49 25b2 − 40b + 16 4x2 − 6x + 9

OpenStudy (anonymous):

The first term has to be a perfect square, the last term has to be a perfect square, and the middle term has to be twice the product of the first and last terms square root

OpenStudy (anonymous):

Which of the polynomials are of the form a^2 -2ab + b^2?

OpenStudy (anonymous):

\[\sqrt{36}=6, \sqrt{9}=3, 3\times2=6\] so its wrong \[\sqrt{4}=2,\sqrt{49}=7, 7\times2=14\] so its wrong \[\sqrt{25}=5,\sqrt{16}=4, 4\times2=8\] so its wrong \[\sqrt{4}=2,\sqrt{9}=3,3\times2=6\] so its right The last one is the perfect square trinomial

OpenStudy (anonymous):

thanks Ashley